OpenStudy (anonymous):

The heat of combustion is 22.7kJ/g which of the following is the theoretical mass of water that could be heated from 25 degrees to 53 degrees using the heat produced by the complete combustion of 0.5g of methanol a) 97g b)0.097g c)194g d)0.194g

OpenStudy (callisto):

Heat released by complete combustion of 0.5g of methanol = 22.7x0.5 = 11.35kJ By E=mcΔT 11.35 = m (4.1813) (53-25) 11.35 = m(4.1813)(28) 11.35 = 117.0764m m = 0.097g